# Universal thermodynamic topological classes of black holes in perfect fluid dark matter background

### v1

@@ Abstract
In this paper, we study the universal thermodynamic topological classes of a family of black holes in a perfect fluid dark matter (PFDM) background. Recent research on black hole thermodynamics suggests that all black holes can be classified into four universal thermodynamic classes, denoted by $W^{1-}$, $W^{0+}$, $W^{0-}$, and $W^{1+}$. Our study reveals that the Schwarzschild black hole in PFDM belongs to the $W^{1-}$ class, and independence of black hole size thermodynamically unstable at both low- and high-temperature limits. The Reissner–Nordström, Kerr, and Kerr–Newman black holes in the PFDM background belong to the same universal thermodynamic class, $W^{0+}$, which represents small, stable black holes and large, unstable black holes at low-temperature limits, whereas no black hole state exists at high temperatures. The AdS black holes behave differently compared to their counterparts in PFDM. The Schwarzschild–AdS black hole belongs to the $W^{0-}$ class, indicating no black hole state at low temperatures, but small, unstable and large, stable black hole states at high temperatures. Furthermore, the Kerr–AdS black hole belongs to the $W^{1+}$ class, characterized by small, stable black holes at low temperatures, large, stable black holes at high temperatures, and unstable, intermediate-sized black holes at both low and high temperatures. These findings uncover the universal topological classifications underlying black hole thermodynamics, offering profound insights into the fundamental principles of quantum gravity.
@@ I Introduction
The prediction of black holes by the theory of General Relativity has been widely tested experimentally, more notably via observational tools of Event Horizon Telescope (EHT) to observe shadows of supermassive black holes eht1 ; eht2 , as well as signature of gravitational wave generated via merger of black holes gw . These observations suggests that black holes are surrounded by magnetic fields and plasma, and actively accreting dust and plasma. One of the earliest proposals for the formation of black holes is through the collapse of a supermassive star (or a spherically symmetric gravitational collapse of pressureless dust) during the later stages of the star’s cycle os ; penrose . However, black holes can also form by the merger of highly dense stars such as neutron stars. Using observed candidates of black holes, the theories of gravity can be tested as well as various black hole solutions can be confronted with observations. From theoretical perspective, novel features of black holes have been already studied. The profound connection between black holes and thermodynamics can be traced back to the groundbreaking work of Hawking hawkingbh . Another important discovery is the interesting connection between the Einstein field equations, the fluid dynamics and thermodynamics pady . In literature, charged and spinning black holes have been tested for stability under accretion of charged or spinning particles to validate the weak cosmic censorship conjecture as well sanjar ; sanjar1 .
Sometime ago, a novel approach for investigating the thermodynamic aspects of black holes has been proposed in Wei:2022dzw . They performed a comprehensive examination of black holes by studying the topological classes of black hole thermodynamics by making use of the generalized off-shell free energy setup. As a result, the manifold of black hole solutions is categorized into three distinct topological classes, their differentiation arising from diverse topological numbers. Within this manuscript, our focus is directed toward an examination of the thermodynamic topological categories pertaining to both stationary and rotating black holes immersed in the framework of PFDM. In literature, several physical aspects of black holes in PFDM background have been investigated including black hole shadows, quasinormal modes, and deflection angles Refs.pfdm0 ; pfdm1 ; pfdm2 ; pfdm3 ; pfdm4 ; pfdm5 ; rizwanatal . The topological approach to study the thermodynamics was used in recent works rtop1 ; rtop2PFDM ; top1 ; top2 ; top3 ; top4 ; top5 ; top6 and, in particular, it was extended to rotating black holes in Ref. Wu:2022whe . In what follows, we shall sketch a few concepts and provide an overview of the topological methodology introduced in Wei:2022dzw . First of all, one has to use the notion of the generalized off-shell Helmholtz free energy, given by
@@
where $M$ and $S$ represent the mass and entropy, respectively, and $\tau$ is the inverse temperature parameter related to the cavity Wei:2022dzw . The generalized Helmholtz free energy exhibits only on-shell properties and reduces to the standard Helmholtz free energy $F=M-TS$ of the black hole when $\tau=\beta=1/T$ rtop3unipaper ; rtop4unitopo . Furthermore, as an integral component of this framework, a core vector $\phi$ which is gradient of the function and can be introduced, as below Wei:2022dzw
@@
~$\tilde{\mathcal{F}}=\mathcal{F}+\frac{1}{\sin\Theta},$ (3)
@@
where $w_{i}$ represents the winding number corresponding to the $i$-th zero point and characterizes the thermodynamical topological class of the black hole. Recent studies suggest that all black holes can be universally classified into four thermodynamical classes: $W^{1-},~{}W^{0+},~{}W^{0-}$, and $W^{1+}$, based on the stable and unstable black hole states indicated by the winding numbers of the field $\phi$ rtop3unipaper ; rtop4unitopo . This topological classification may provide insights into thermodynamic properties, contributing to a better understanding of black hole physics.
This paper is organized as follows: In the subsequent section, we study the topological the universal topological classifications concerning family of static and stationary black holes within the context of PFDM. Moving forward, in sections II our focus shifts to an examination of the topological classes associated Schwarzschild black holes in the presence of PFDM. We extend our analysis in section III to encompass the Reissner–Nordström black hole existing within the PFDM framework. Additionally, in section IV, we discuss the topological classes of Kerr, and Kerr–Newman black. In section V and VI, we discuss the how the Schwarszchild-Ads and Kerr-Ads black holes can be classify thermodynamically. Our paper culminates with a presentation of conclusions, outlined in section VI. We shall adopt units such that $c=1=G.$
@@ II Schwarzschild Black Hole in a Dark Matter Background
Note that, in the absence of the PFDM, that is, for $\alpha=0$, the line element and the thermodynamical quantities reduce to that of the Schwarzschild black hole. For the Schwarzschild black hole in PFDM background the asymptotic behaviour of the inverse temperature parameter $\beta$ exhibits the following limits:
$\beta(r_{m})=0\quad\text{and}\quad\beta(\infty)=\infty,$ (8)
@@
Figure 3: Contours $\Phi_{1}$ represent the change in the components of the vector field $\phi$ as the contour $C_{1}$ in Fig. 1 is traversed for a Schwarzschild black hole in PFDM. The black dot indicates the zero point of the vector field, and the arrows show the sense of rotation that the vector $\phi$ undergoes while traversing the contour $C_{1}$. The changes in the components of the field along this closed loop cause $\Phi_{1}$ to rotate in a clockwise direction, corresponding to a winding number of $-1$.
@@
Now, we study the behavior of the vector field $\phi$ (plotted in Fig. 1) at the boundaries corresponding to Eq. (8). The overall boundary is described by a contour $C=I_{1}\cup I_{2}\cup I_{3}\cup I_{4}$, where the segments $I_{i}$ and the contour $C$ are shown in Fig. 2, encompassing all regions of the parameters. The setup of $\phi$ indicates that it is orthogonal to $I_{1}$ and $I_{3}$ rtop1 , so the key asymptotic behavior occurs along $I_{2}$ and $I_{4}$. As $r_{h}\to r_{m}$, the vector field plot of $\phi$ suggests that the direction of $\phi$ is rightward with an inclination depending on $\Theta$, whereas when $r_{h}\to\infty$, it is leftward directed (see Fig. 1 and 2) suggesting it thermodynamical class.
@@ III Reissner–Nordström black hole in dark matter background
Now we focus on understanding the importance of the electric charge concerning the topological numbers of thermodynamics. To find the spacetime geometry, one can start from the action for the gravity theory minimally coupled with gauge field in PFDM reAdS as Das:2020yxw
@@
Above $T_{\mu\nu}^{M}$, represents the energy-momentum tensor for ordinary matter while $T_{\mu\nu}^{DM}$ denotes the energy-momentum tensor for PFDM
@@
where $r_{m}=r_{e}$ with $r_{e}$ being the radius of of extremal black hole given by Eq. (19).
@@ IV Kerr-Newman Black Hole in Dark Matter Background
Here, the thermodynamic quantities necessary for the topological classification of the Kerr-Newman black hole in PFDM can be presented as follows MJP
@@
and the the components of the vector field $\phi$ can be evaluates as
@@
The components of the vector field $\phi$ given by Eq. (34) for the Kerr-Newman black hole in the PFDM background have the same form as those obtained for the Reissner–Nordström black hole in the PFDM background, as given by Eq. (23), but with the difference that $Q^{2}\to a^{2}+Q^{2}$. The graphical study of the Kerr-Newman black hole in the PFDM background, presented in Figs. 6 and 7, is also the same. Therefore, we can conclude that the Kerr-Newman (and similarly, Kerr) black holes in the PFDM background belong to the same universal thermodynamical class, $W^{0+}$, as the Reissner–Nordström black hole in PFDM, and the thermodynamical states of these black holes are summarized in Table 1 and 3.
@@
Schwarzschild black hole in PFDM $\leftarrow$ $\uparrow$ $\rightarrow$ $\uparrow$ -1
@@ V Schwarzschild-AdS Black Hole in Dark Matter Background
In this section, we will discuss the topological classes of the Schwarzschild-AdS black hole in PFDM background. The line element of the Schwarzschild-AdS black hole in PFDM background can be written as KerrDM
@@
Here, $\Lambda$ represents the cosmological constant. For all values of the black hole parameters, the Schwarzschild-AdS black hole in PFDM background has a single horizon, represented by $r_{h}$, and in this case, the minimal radius $r_{m}$ of the black hole is zero. The mass, entropy and the Hawking temperature of the Schwarzschild-AdS black hole in PFDM background is given by MJP
@@
These thermodynamic quantities lead to the off-shell free energy, and the corresponding components of the $n$-vector field $\phi$ for the Schwarzschild-AdS black hole are given as follows:
$\mathcal{F}=\frac{1}{6}\left[3r_{h}+8\pi Pr_{h}^{3}+3\alpha\ln\left(\frac{r_{h}}{|\alpha|}\right)\right]-\frac{\pi r_{h}^{2}}{\tau}$ (41)
@@
where $P=-\Lambda/8\pi$ representing the pressure parameter. The vector field $\phi$ for the Schwarzschild-AdS black hole in the PFDM background, is plotted in Fig. 8. The graph shows that, unlike the Schwarzschild black hole in PFDM, the Schwarzschild-AdS black hole in PFDM has two zero points, represented by black dots. The asymptotic behavior of the vector field $\phi$ at the boundaries $I_{i}$ is shown in Table 2. It is observed that as $r_{h}\to r_{m}$ or $r\to\infty$, the vector field plot of $\phi$ indicates a rightward direction, suggesting its universal thermodynamical class.
For the Schwarzschild-AdS black hole in the PFDM background, the contours $\Phi_{i}$ that map the change in the components $(\phi^{r_{h}},\phi^{\Theta})$ of the vector field $\phi$ as the contours $C_{i}$ in Fig. 8 are traversed are plotted in Fig. 9. The zero point of the field $\phi$ is at the origin, and contours with negative winding number are mapped to clockwise loop, whereas with a positive winding number is mapped to counterclockwise loop. A annihilation point is found at $\beta=\beta_{c}$ (which depends on the parametric values of the black hole parameters). For $\beta<\beta_{c}$, there are two black hole states: small black holes with winding number $-1$, and large black holes with winding number $1$, representing thermodynamically unstable and stable black holes, respectively. However, for $\beta_{c}<\beta$, no black hole state is present. The topological classes remain unaffected regardless of the values of $\beta$ or other black hole parameters.
@@ VI Kerr-AdS Black Hole in Dark Matter Background
Here $m$ and $a$ are the mass and angular momentum per unit mass parameters of the black hole and $\alpha$ and $\Lambda$ are the PFDM parameter and the cosmological constant, respectively.
@@
Table 3: Universal thermodynamic topological classifications of family of the black holes in PFDM background and their thermodynamical properties. DG, AP, and GP are for degenerate point, annihilation point, and generation point, respectively.
@@
The mass and other thermodynamical quantities of Kerr-AdS black hole PFDM background are given by given by MJP
@@
Using these thermodynamical quantities generalized off-shell free energy of the system can be obtained as
@@
The vector field $\phi$ for the Kerr-AdS black hole in the PFDM background is plotted in Fig. 10, which shows that there are three zero points of the vector field, denoted by black dots. These zero points divide the black hole into a small, intermediate, and large black hole, with one generation point and one annihilation point. Furthermore, the asymptotic behavior of the vector field at the boundaries $l_{i}$ is shown in Table 2. The direction of the vector field at the boundary $I_{1}$ is rightward, while at the boundary $I_{3}$, it is leftward, suggesting its universal class.
For the Kerr-AdS black hole in PFDM, we have two critical values, $\tau_{a}$ and $\tau_{b}$. For $\tau<\tau_{a}$ or $\tau_{b}<\tau$, there is only black hole state with winding number $1$ representing large and small size black hole respectively . However, for $\tau_{a}<\tau<\tau_{b}$, three black hole states exist: small, intermediate, and large black holes. The contours $\Phi_{i}$, which map the changes in the components $\left(\phi^{r_{h}},\phi^{\theta}\right)$ of the vector field $\phi$ as the contours $C_{i}$ shown in Fig. 10 are traversed, are plotted in Fig. 11. These contours yield the winding numbers of the zero points as $1$, $-1$, and $1$, respectively. Thus, the small and large black holes have positive heat capacity and are stable, whereas the intermediate black holes have negative heat capacity and are unstable.

### v2

# Universal Thermodynamic Topological Classes of Black Holes in a Perfect Fluid Dark Matter Background

@@ Abstract
In this paper, we study the universal thermodynamic topological classes of a family of black holes in a perfect fluid dark matter (PFDM) background. Recent research on black hole thermodynamics suggests that all black holes can be cast into four universal thermodynamic classes, denoted by $W^{1-}$, $W^{0+}$, $W^{0-}$, and $W^{1+}$. Our study reveals that the Schwarzschild black hole in PFDM belongs to the $W^{1-}$ class, and the independence of black hole size is thermodynamically unstable in both the low- and high-temperature limits. The Reissner–Nordström, Kerr, and Kerr–Newman black holes in the PFDM background belong to the same universal thermodynamic class, $W^{0+}$, which represents small-stable black holes and large-unstable black holes at low-temperature limits, whereas no black hole state exists at high temperatures. The AdS black holes behave differently when compared to their counterparts in PFDM. The Schwarzschild–AdS black hole belongs to the $W^{0-}$ class, indicating that no black hole state exists at low temperatures, but small-unstable and large-stable black hole states exist at high temperatures. Furthermore, the Kerr–AdS black hole belongs to the $W^{1+}$ class, characterized by small-stable black holes at low temperatures, large-stable black holes at high temperatures, and unstable intermediate-sized black holes at both low and high temperatures. Thus, the presence of PFDM does not affect the stability of the black hole. These findings uncover the universal topological classifications underlying black hole thermodynamics, offering profound insights into the fundamental principles of quantum gravity.
@@ I Introduction
The prediction of black holes by the theory of General Relativity has been widely tested experimentally, more notably via the Event Horizon Telescope (EHT) to observe shadows of supermassive black holes eht1 ; eht2 , as well as signature of gravitational wave generated via merger of black holes via LIGO/VIRGO observatories gw . The astronomical observations suggest that black holes are surrounded by magnetic fields and plasma and actively accreting dust and plasma. One of the earliest proposals for the formation of black holes is through the collapse of a supermassive star (or a spherically symmetric gravitational collapse of pressureless dust) during the final stages of the star’s evolution os ; penrose . However, black holes can also be formed by the merger of highly dense stars such as neutron stars ns . Using observed candidates of black holes, the theories of modified gravity can be tested by confronting various black hole solutions in modified gravity theories with the astronomical observations. From theoretical perspective, novel features of black holes have been already studied. The profound connection between black holes and thermodynamics can be traced back to the groundbreaking work of Hawking hawkingbh . Another important discovery is the interesting connection between the Einstein field equations, the fluid dynamics and thermodynamics pady . In literature, charged and spinning black holes have been tested for stability under accretion of charged or spinning particles to validate the weak cosmic censorship conjecture as well sanjar ; sanjar1 .
A novel approach for investigating the thermodynamic aspects of black holes has been proposed in Wei:2022dzw , where the authors performed a comprehensive examination of black holes by studying the topological classes of black hole thermodynamics using the generalized off-shell free energy setup. As a result, the manifold of black hole solutions is categorized into three distinct topological classes, their differentiation arising from diverse topological numbers. Here, our focus is the examination of the thermodynamic topological categories pertaining to both stationary and axi-symmetric black holes immersed in PFDM. In literature, several physical aspects of black holes in PFDM background have been investigated including black hole shadows, quasinormal modes, and deflection angles pfdm0 ; pfdm1 ; pfdm2 ; pfdm3 ; pfdm4 ; pfdm5 ; rizwanatal . The topological approach to study the black hole thermodynamics was used in recent works rtop1 ; rtop2PFDM ; top1 ; top2 ; top3 ; top4 ; top5 ; top6 and, in particular, extended to rotating black holes in Wu:2022whe . In what follows, we shall sketch a few concepts and provide an overview of the topological methodology introduced in Wei:2022dzw . To proceed, the notion of the generalized off-shell Helmholtz free energy is given by
@@
where $M$ and $S$ represent the mass and entropy, respectively, and $\tau$ is the inverse temperature parameter related to the cavity Wei:2022dzw . The generalized Helmholtz free energy exhibits only on-shell properties and reduces to the standard Helmholtz free energy $F=M-TS$ of the black hole when $\tau=\beta=1/T$ rtop3unipaper ; rtop4unitopo . Furthermore, as an integral component of this framework, a vector field $\phi$ which is gradient of the function can be introduced as Wei:2022dzw
@@
~$\tilde{\mathcal{F}}=\mathcal{F}+\frac{1}{\sin\Theta}.$ (3)
@@
where $w_{i}$ represents the winding number corresponding to the ith zero point and characterizes the thermodynamical topological class of the black hole. Recent studies suggest that all black holes can be universally classified into four thermodynamical classes: $W^{1-},~{}W^{0+},~{}W^{0-}$, and $W^{1+}$, based on the stable and unstable black hole states indicated by the winding numbers of the vector field $\phi$ rtop3unipaper ; rtop4unitopo . This topological classification may provide insights into thermodynamic properties, contributing to a better understanding of black hole physics.
This paper is organized as follows: In sections II we do an examination of the topological classes associated Schwarzschild black holes in the presence of PFDM. We extend the analysis in section III to encompass the Reissner–Nordström black hole existing within the PFDM framework. In section IV, the topological classes of Kerr, and Kerr–Newman black are discussed. In section V and VI, we discuss how can the Schwarszchild-AdS and Kerr-AdS black holes be classified thermodynamically. Our paper culminates with a presentation of conclusions, outlined in section VII. We shall adopt units such that $c=1=G.$
@@ II Schwarzschild Black Hole in a Dark Matter Background
Note that, in the absence of the PFDM, $\alpha=0$, the line element and the thermodynamical quantities reduce to that of the Schwarzschild black hole. For the Schwarzschild black hole in PFDM background the asymptotic behaviour of the inverse temperature parameter $\beta$ admits the following limits:
$\beta(r_{m})=0,\quad\text{and}\quad\beta(\infty)=\infty,$ (8)
@@
Figure 3: Contour $\Phi_{1}$ represents the change in the components of the vector field $\phi$ as the contour $C_{1}$ in Fig. 1 is traversed for a Schwarzschild black hole in PFDM. The black dot indicates the zero point of the vector field, and the arrows show the sense of rotation that the vector $\phi$ undergoes while traversing the contour $C_{1}$. The changes in the components of the field along this closed loop cause $\Phi_{1}$ to rotate in a clockwise direction, corresponding to a winding number of $-1$.
@@
Now, we study the behavior of the vector field $\phi$ (plotted in Fig. 1) at the boundaries corresponding to Eq. (8). The overall boundary is described by a contour $C=I_{1}\cup I_{2}\cup I_{3}\cup I_{4}$, where the segments $I_{i}$ and the contour $C$ are shown in Fig. 2, encompassing all regions of the parameters. The setup of $\phi$ indicates that it is orthogonal to $I_{1}$ and $I_{3}$ rtop1 , therefore the key asymptotic behavior occurs along $I_{2}$ and $I_{4}$. As $r_{h}\to r_{m}$, the vector field plot of $\phi$ suggests that the direction of $\phi$ is rightward with an inclination depending on $\Theta$, whereas when $r_{h}\to\infty$, it is leftward directed (see Fig. 1 and 2) suggesting it thermodynamical class.
@@ III Reissner–Nordström black hole in dark matter background
Now we investigate the impact of the electric charge in the topological numbers of thermodynamics. To find the spacetime geometry, one can proceed using the action for the gravity theory minimally coupled with gauge field in PFDM, which is read as Das:2020yxw
@@
Above $T_{\mu\nu}^{M}$ and $T_{\mu\nu}^{DM}$ represent the energy-momentum tensor for ordinary matter and PFDM respectively,
@@
where $r_{m}=r_{e}$ with $r_{e}$ being the radius of extremal black hole given by Eq. (19).
@@ IV Kerr-Newman Black Hole in Dark Matter Background
Here, the thermodynamic quantities to investigate the topological classification of the Kerr-Newman black hole in PFDM can be presented as MJP
@@
and the components of the vector field $\phi$ are evaluated as
@@
The components of the vector field $\phi$ given by Eq. (34) for the Kerr-Newman black hole in the PFDM background have the same form as those obtained for the Reissner–Nordström black hole in the PFDM background, as given by Eq. (23), provided $Q^{2}\to a^{2}+Q^{2}$. The graphical study of the Kerr-Newman black hole in the PFDM background, presented in Figs. 6 and 7, is also the same. Therefore, we can conclude that the Kerr-Newman (and similarly, Kerr) black holes in the PFDM background belong to the same universal thermodynamical class, $W^{0+}$, as the Reissner–Nordström black hole in PFDM, and the thermodynamical states of these black holes are summarized in Table 1 and 3.
@@
Schwarzschild black hole in PFDM $\leftarrow$ $\uparrow$ $\rightarrow$ $\downarrow$ -1
@@ V Schwarzschild-AdS Black Hole in Dark Matter Background
Now we discuss the topological classes of the Schwarzschild-AdS black hole in PFDM background whose line element can be written as KerrDM
@@
Here $\Lambda<0$ denotes the cosmological constant. For all values of the black hole parameters, the Schwarzschild-AdS black hole in PFDM background has a single horizon, represented by $r_{h}$, and in this case, the minimal radius $r_{m}$ of the black hole is zero. The mass, entropy and the Hawking temperature of the Schwarzschild-AdS black hole in PFDM background is given by MJP
@@
These thermodynamic quantities lead to the off-shell free energy, and the corresponding components of the $n$-vector field $\phi$ for the Schwarzschild-AdS black hole have the form:
$\mathcal{F}=\frac{1}{6}\left[3r_{h}+8\pi Pr_{h}^{3}+3\alpha\ln\left(\frac{r_{h}}{|\alpha|}\right)\right]-\frac{\pi r_{h}^{2}}{\tau},$ (41)
@@
where $P=-\Lambda/8\pi$ representing the pressure parameter. The vector field $\phi$ for the Schwarzschild-AdS black hole in the PFDM background, is shown in Fig. 8. The figure shows that, unlike the Schwarzschild black hole in PFDM, the Schwarzschild-AdS black hole in PFDM has two zero points, represented by black dots. The asymptotic behavior of the vector field $\phi$ at the boundaries $I_{i}$ is shown in Table 2. It is observed that as $r_{h}\to r_{m}$ or if $r\to\infty$, the vector field plot of $\phi$ indicates a rightward direction, suggesting its universal thermodynamical class.
For the Schwarzschild-AdS black hole in the PFDM background, the contours $\Phi_{i}$ that map the change in the components $(\phi^{r_{h}},\phi^{\Theta})$ of the vector field $\phi$ as the contours $C_{i}$ in Fig. 8 are traversed are plotted in Fig. 9. The zero point of the field $\phi$ is at the origin, and contours with negative winding number are mapped to clockwise loop, whereas with a positive winding number is mapped to counterclockwise loop. An annihilation point exists at $\beta=\beta_{c}$ (which depends on the parametric values of the black hole parameters). For $\beta<\beta_{c}$, there are two black hole states: small black holes with winding number $-1$, and large black holes with winding number $1$, representing thermodynamically unstable and stable black holes, respectively. However, for $\beta_{c}<\beta$, no black hole state is present. The topological classes remain unaffected regardless of the values of $\beta$ or other black hole parameters.
@@ VI Kerr-AdS Black Hole in Dark Matter Background
Here $m$ and $a$ denote the mass and angular momentum per unit mass of the black hole while $\alpha$ and $\Lambda$ are the PFDM parameter and the cosmological constant, respectively.
@@
Table 3: Universal thermodynamic topological classifications of family of the black holes in PFDM background and their thermodynamical properties. DP, AP and GP are used to represent degenerate point, annihilation point, and generation point, respectively.
@@
The mass and other thermodynamical quantities of Kerr-AdS black hole PFDM background are given by MJP
@@
Using these thermodynamical quantities, the generalized off-shell free energy of the system can be obtained as
@@
The vector field $\phi$ for the Kerr-AdS black hole in the PFDM background is plotted in Fig. 10, which shows that there are three zero points of the vector field, denoted by black dots. These zero points divide the black hole configuration into a small, intermediate, and large black hole type, with one generation point and one annihilation point. Furthermore, the asymptotic behavior of the vector field at the boundaries $l_{i}$ is shown in Table 2. The direction of the vector field at the boundary $I_{1}$ is rightward, while at the boundary $I_{3}$, it is leftward, suggesting its universal class.
For the Kerr-AdS black hole in PFDM, we have two critical values, $\tau_{a}$ and $\tau_{b}$. For $\tau<\tau_{a}$ or $\tau_{b}<\tau$, there is only black hole state with winding number $1$ representing large and small size black hole respectively . However, for $\tau_{a}<\tau<\tau_{b}$, three black hole states exist: small, intermediate, and large black holes. The contours $\Phi_{i}$, which map the changes in the components $\left(\phi^{r_{h}},\phi^{\theta}\right)$ of the vector field $\phi$ as the contours $C_{i}$ shown in Fig. 10 are traversed, are plotted in Fig. 11. These contours yield the winding numbers of the zero points as $1$, $-1$, and $1$, respectively. Thus, the small and large black holes have positive heat capacity and hence are stable, whereas the intermediate black holes have negative heat capacity and are consequently unstable.